A satellite moves in a circular orbit at a constant speed around the Earth. Which of the following statements is true?
(a) No force acts on the satellite
(b) The satellite moves at a constant speed and hence doesn't accelerate.
(c) The satellite has an acceleration directed away from the Earth.
(d) The satellite has an acceleration directed toward the Earth.
1
Expert's answer
2021-02-18T18:43:08-0500

A circular motion having a constant speed never means that it does not accelerate because in circular motion, direction of motion always changes. So, velocity of satellite is varying every moment which requires an acceleration known as centripetal acceleration and in other words , require a centripetal force which act on the particle and is responsible for circular motion.

The magnitude of acceleration and force are given by ,

"a_c=\frac{v^2}{r}\ ; F_c=\frac{mv^2}{r}" directed towards the center of circle.

Hence , Option "(d)" is correct ,The satellite has an acceleration directed toward the Earth.